# Making school choice lotteries transparent

### v2

@@ 3.1 Experimental Design
The experiment was conducted in May 2025 at the Economics Experimental Lab of Nanjing Audit University with a total of 240 university students, using the software z-Tree (Fischbacher, 2007). We ran four sessions for each treatment. Each session consisted of 12 participants who interacted for a total of 20 rounds. In each round, participants were randomly assigned to two 6-person groups. Within each group, each participant was randomly assigned a student ID (ranging from 1 to 6), a lottery number (also ranging from 1 to 6), and payoffs associated with attending each of the three schools; all of these values were distinct in each round. Specifically, we generated a sequence of the 5-tuples, consisting of the student ID, lottery number, and three payoffs, at the group-round level. We applied the sequence to each session to alleviate concerns that treatment effects might arise solely from differences in random number generation. After each round, participants received feedback about whether and to which school they were admitted and their round payoff. At the end of the session, one round was privately and randomly chosen for each participant to determine their final payoff. The experimental instructions are available in Online Appendix B.
@@ 3.2 Experimental Results
Table 4 reports the observed match rate for each treatment. The Reveal policy yields a significantly higher match rate than the Cover policy, irrespective of the presence of neighborhood schools. In the absence of neighborhood schools, the average match rate under the Reveal policy is 96.25%, which is significantly higher than 67.81% under the Cover policy ($p$ = 0.029, Wilcoxon rank-sum test with each session treated as an independent observation; same below). [^17] With 4 sessions per treatment, 0.029 is the lowest possible p-value when conducting this test. Similarly, when neighborhood schools are available, the match rates are 93.54% and 87.08% under the two reveal policies (NS_RevealMore and NS_Reveal, respectively), compared to 77.92% under the Cover policy ($p$ = 0.029). Importantly, the additional information about the lottery numbers of students with neighborhood schools provided in NS_RevealMore significantly improves the match rate relative to the standard Reveal policy in NS_Reveal ($p$ = 0.029). [^18] Online Appendix Figure A1 illustrates the match rate across rounds for each treatment, suggesting that the observed differences persist throughout the experiment. These findings are consistent with Hypothesis 1.
@@
To summarize participants’ adherence to equilibrium strategies across treatments, we construct a statistic termed average absolute deviation (AAD), which calculates the treatment-level frequency of deviating from the equilibrium strategies summarized in Table 2. 2(b) displays the results, showing that the Reveal policy substantially reduces AAD irrespective of the presence of neighborhood schools. In the absence of neighborhood schools, the Reveal policy reduces AAD from 0.42 to 0.04 ($p$ = 0.029), meaning that only 4% of decisions deviate from equilibrium strategies. When neighborhood schools are available, the Reveal policy reduces AAD from 0.39 to 0.21, and the RevealMore policy further reduces it to 0.07 ($p$ = 0.029 in all treatment comparisons). 2(b) further shows that treatment differences persist across rounds, although AAD generally decreases over time, reflecting learning effects; this reduction is especially pronounced in NS_Reveal. [^19] Online Appendix Table A1 reports complementary regression analyses confirming that treatment differences are significant both across all rounds and during the last 5 rounds, with a noticeable decline in suboptimal decisions in NS_Reveal during the last 5 rounds.
@@
Note that in NS_Reveal, no-neighborhood students’ strategies are “non-monotonic” in their lottery numbers. Specifically, their equilibrium strategy is to report $s_{1}$ when the lottery number is 5, which is unintuitive. Indeed, experimental data indicate that participants deviated from this strategy over 95% of the time under this circumstance (see Table 5). Therefore, we performed a robustness check by excluding observations for no-neighborhood students with lottery number 5 across all treatments (to ensure fair comparisons). AAD in NS_Reveal decreases to 0.14, which remains significantly higher than that in NS_RevealMore (pp=0.029). However, Online Appendix 2(b) suggests that during the last 5 rounds, AAD levels between NS_Reveal and NS_RevealMore become nearly indistinguishable, a finding corroborated by complementary regression analyses reported in Online Appendix Table A2. This finding suggests that the efficiency gain associated with providing the additional lottery information tends to diminish as participants gain experience. This is nonetheless consistent with our earlier discussion at the end of Subsection 2.2, which posits that increased market stability may reduce the necessity for full revelation of lottery information. Overall, the analysis of AAD offers initial empirical support for Hypothesis 3.
@@
Table 5 shows that, in the absence of neighborhood schools, the Reveal policy significantly increases the likelihood that students with various lottery numbers adhere to equilibrium strategies. When neighborhood schools are available, the Reveal policy mainly benefits $s_{3}$-neighbor and no-neighborhood students by increasing their conformity to equilibrium strategies. [^20] The Reveal policy also slightly reduces the overall frequency of deviation for $s_{2}$-neighbor students from 8.75% to 6.25%. However, $s_{2}$-neighbor students with lottery number 2 tend to make more mistakes than those with other lottery numbers, occasionally reporting $s_{1}$ despite an equilibrium strategy to always report $s_{2}$. For example, $s_{3}$-neighbor students deviate only 15% of the time, which is substantially lower than 68.13% in NS_Cover. [^21] In previous literature, this behavioral pattern of $s_{3}$-neighbor students observed in NS_Cover is referred to as neighborhood school bias (Calsamiglia, Haeringer and Klijn, 2010). However, given that theoretical predictions vary across policies, we refrain from asserting that the Reveal policy directly mitigates this bias here. Notably, the efficacy of the Reveal policy is not uniform across lottery numbers; students receiving either the best or worst lottery numbers deviate much less frequently than those with intermediate lottery numbers. As previously noted, the equilibrium strategy for no-neighborhood schools with lottery number 5 is unintuitive, causing almost all these students’ decisions to deviate from the equilibrium strategy. However, even after excluding these observations, the tendency of $s_{3}$-neighbor and no-neighborhood students with intermediate lottery numbers making more mistakes remains persistent. To further investigate the sources of these mistakes, Table 6 presents the frequencies of all possible submission strategies by students’ roles and utility types. Compared to theoretical predictions summarized in Table 3, $s_{3}$-neighbor and no-neighborhood students tend to err on the side of being too conservative in their submission strategies under the Cover policy. Although the Reveal policy mitigates this tendency, particularly among type-$\mathbf{v}$ students, it does not entirely eliminate it. This behavioral pattern suggests that $s_{3}$-neighbor and no-neighborhood students with intermediate lottery numbers face more complex strategic challenges, which the Reveal policy partially alleviates, relative to $s_{1}$-neighbor, $s_{2}$-neighbor and students with either the best or worst lottery numbers.
@@
Finally, panel (3) of Table 5 shows that the RevealMore policy further enhances adherence to equilibrium strategies among $s_{3}$-neighbor and no-neighborhood students, notwithstanding the persistence of relatively higher deviation rates among those with intermediate lottery numbers. For example, $s_{3}$-neighbor students deviate only 6.25% of the time; even those with lottery number 3 err in just 14.29% of cases, substantially lower than 46.43% in NS_Reveal. Online Appendix Table A3 reports more detailed analyses, allowing decisions to depend on both own lottery number and lottery numbers of neighborhood students. We observe that except in a few instances with limited observations (i.e., no more than 6), the frequencies of deviation generally remain below 20%. One notable exception occurs when no-neighborhood students with lottery number 4 observe that both $s_{1}$-neighbor’s and $s_{2}$-neighbor’s lottery numbers exceed 2. In this case, although they should have reported $s_{3}$, they mistakenly report $s_{2}$ in 35% of the time. This discrepancy likely reflects the strategic complexity inherent in the situation: despite holding a better-than-average lottery number of 3, which intuitively suggests a favorable chance of admission to $s_{2}$, the equilibrium strategy indicates otherwise.
@@ 3.3 A Robustness Experiment in More Complex Environments
Our main experiment examines scenarios in which students are permitted to report only one school in their preference list. The matching market is relatively small and the cardinal preferences are somewhat arbitrary. However, this setup is essential for making precise theoretical predictions, but perhaps at the cost of realism. To partially address this limitation, the robustness experiment considers a larger matching market with neighborhood schools (16 students and 4 schools), allowing students to report up to two schools. All students’ ordinal preferences are identical, but their cardinal preferences incorporate multiple random components. The principal findings are consistent across both experiments: the Reveal policy improves the overall match rate and market efficiency compared to the Cover policy, and students generally make more informed decisions based on observed lottery numbers. A detailed report on the robustness experiment is provided in Online Appendix C.
@@ 4 Conclusion
Lotteries are widely used in school choice systems to break priority ties. This paper advocates for a policy that reveals lottery outcomes to students before preference submission, contrasting it with the prevailing practice of withholding such information. Our theoretical analysis shows that when students face constraints on the length of their preference lists, which is a common feature in practice, the reveal policy improves students’ decision-making, enabling them to coordinate their choices more effectively. This, in turn, enhances matching outcomes and increases overall welfare. Complementary lab experiments support these findings, showing that the reveal policy leads to higher match rates , higher student payoffs, and better alignment between reported preferences and achievable options.
@@
Appendix A Additional Figures and Tables
@@
Appendix B Experimental Instructions
@@
Appendix C Robustness Experiment
C.1 Experimental Design
@@
The experiment was conducted at the Nanjing Audit University Economics Experimental Lab with a total of 192 university students, using the software z-Tree (Fischbacher, 2007). We ran six sessions for each treatment. Each session consisted of 16 participants who interacted for a total of 40 rounds, divided into two blocks of 20 rounds each. Within each group, each participant was randomly assigned a student ID (ranging from 1 to 16), a lottery number (also ranging from 1 to 16), and payoffs from attending each of the four schools; all of these values were distinct in each round. Specifically, we generated a sequence of the 6-tuple of ID, lottery number and four payoffs at the group-round level. We applied it to each session to alleviate the concern that any treatment difference is simply due to the different random numbers. After every round, all participants received feedback about whether and which school they were admitted to and their round payoff. At the end of the session, one round from each block was privately and randomly chosen for each participant as her payoff for that block, and the participant’s total payoff was the sum of the payoffs from the two blocks. The experimental instructions are reproduced in the online appendix.
@@ C.2 Experimental Results
Figure C1 shows the match rate across rounds under each condition. Consistent with our hypothesis, the reveal policy results in a significantly higher match rate than the cover policy, irrespective of the ROL length. When the ROL length is one, the average match rate is 89.1% under the reveal policy, which is significantly higher than 77.8% under the cover policy ($p$ = 0.004, Wilcoxon rank-sum test with each session considered as an independent observation). Similarly, when the ROL length is two, the rate is 98.5% under the reveal policy compared to 85.9% under the cover policy ($p$ = 0.003).
Figure C1: Match rate over round
@@
The reveal policy is expected to assist students in making more rational decisions and thereby reduce their propensity to play safe by frequently reporting neighborhood schools. Table C1 presents the neighborhood school report rate under each condition, focusing on students who have neighborhood schools. The table also presents a breakdown of this statistic by each of the four lottery categories. For instance, lottery (1, 4) refers to the four students whose lottery numbers are 1, 2, 3, and 4.
@@
Table C1: Neighborhood school report rate and bias
@@
The reveal policy is expected to reduce not only the neighborhood school bias in students’ ROL submissions but also the segregation level in the final admission outcome, measured as the proportion of students assigned to their neighborhood schools. Table C2 in Online Appendix A indicates that the reveal policy results in a significantly lower segregation level compared to the cover policy, but this effect is significant only when the ROL length is one. When the ROL length is two, although the effect aligns with expectations, the reveal policy exhibits a weaker and statistically insignificant effect on the segregation level. This is primarily driven by the luckiest students who could still be admitted to schools that are better than their neighborhood schools when they have more than one school to report. Thus, their neighborhood school bias has a much weaker impact on their final admission outcomes.
Table C2: Segregation (the proportion of subjects assigned to their neighborhood schools, only for those who have a neighborhood school)
@@
Table C3 reports the efficiency, measured by the student’s average payoff, under each condition. Consistent with our hypothesis, the reveal policy results in significantly greater overall efficiency than the cover policy, irrespective of the ROL length. However, the benefits of the reveal policy are not uniformly distributed among students. While students with the best or worst lottery numbers tend to benefit from this policy, students with intermediate lottery numbers tend to suffer. The issue of unevenly distributed benefits is pronounced when the ROL length is one, but it is largely mitigated when the ROL length is two. Intuitively, when the possibility of mismatch is high (i.e., when the ROL length is short), prior knowledge of lottery numbers assists the most fortunate students in gaining admission to better schools while helping the least fortunate students avoid being left unmatched. However, this dynamic tends to disadvantage students in the middle, making them less likely to gain admission to top schools.
Table C3: Efficiency (mean payoff)
@@
Following our review of the aggregate-level results, we turn to individual ROL strategies. Table C4 presents the frequency of each possible strategy under each condition. Under the reveal policy, we also calculate the frequency for each lottery category separately. When the ROL length is one, students tend to report better schools more frequently under the cover policy; in contrast, under the reveal policy, the frequency of reporting each school is relatively uniform. The reason is that students make more informed decisions under the reveal policy by primarily reporting schools where they have a realistic chance of admission: students whose lottery numbers range from 1 to 4 are most likely to report school A; students whose lottery numbers range from 5 to 8 are most likely to report school B; and so on. Similarly, when the ROL length is two, under the cover policy, students tend to report a pair of schools that more frequently include at least one top school (such as A-B and A-C) than two bottom schools (such as C-D); conversely, under the reveal policy, the reported pairs of schools are evenly distributed between top and bottom schools (such as A-B, B-C and C-D). Further, students tend to report schools to which they have realistic chances of admission based on their lottery numbers: students whose lottery numbers range from 1 to 4 are most likely to report the school pair A-B; students whose lottery numbers range from 5 to 8 are most likely to report the school pair B-C; students whose lottery numbers range from 9 to 16 are most likely to report the school pair C-D. [^22] We also calculate the frequency of each possible strategy for students who have a specific neighborhood school (Table C5, Table C6 and Table C7) and for students who do not have any neighborhood school (Table C8) separately. The results generally align with our expectations. For instance, students who have neighborhood school B only report school A or B when the ROL length is one and the school pairs A-B, B-C or B-D when the ROL length is two.
Table C4: Proportion of each possible submission strategy (%)
@@
Table C5: Proportion of each possible submission strategy (%) for students having neighborhood school B
@@
Table C6: Proportion of each possible submission strategy (%) for students having neighborhood school C
@@
Table C7: Proportion of each possible submission strategy (%) for students having neighborhood school D
@@
Table C8: Proportion of each possible submission strategy (%) for non-neighborhood students

### v3

@@ 3.1 Experimental Design
The experiment was conducted in May 2025 at the Economics Experimental Lab of Nanjing Audit University with a total of 240 university students, using the software z-Tree (Fischbacher2007). We ran four sessions for each treatment. Each session consisted of 12 participants who interacted for a total of 20 rounds. In each round, participants were randomly assigned to two 6-person groups. Within each group, each participant was randomly assigned a student ID (ranging from 1 to 6), a lottery number (also ranging from 1 to 6), and payoffs associated with attending each of the three schools; all of these values were distinct in each round. Specifically, we generated a sequence of the 5-tuples, consisting of the student ID, lottery number, and three payoffs, at the group-round level. We applied the sequence to each session to alleviate concerns that treatment effects might arise solely from differences in random number generation. After each round, participants received feedback about whether and to which school they were admitted and their round payoff. At the end of the session, one round was privately and randomly chosen for each participant to determine their final payoff. The experimental instructions are available in Online Appendix A.
@@ 3.2 Experimental Results
Table 4 reports the observed match rate for each treatment. The Reveal policy yields a significantly higher match rate than the Cover policy, irrespective of the presence of neighborhood schools. In the absence of neighborhood schools, the average match rate under the Reveal policy is 96.25%, which is significantly higher than 67.81% under the Cover policy ($p$ = 0.029, Wilcoxon rank-sum test with each session treated as an independent observation; same below). [^18] With 4 sessions per treatment, 0.029 is the lowest possible p-value when conducting this test. Similarly, when neighborhood schools are available, the match rates are 93.54% and 87.08% under the two reveal policies (NS_RevealMore and NS_Reveal, respectively), compared to 77.92% under the Cover policy ($p$ = 0.029). Importantly, the additional information about the lottery numbers of students with neighborhood schools provided in NS_RevealMore significantly improves the match rate relative to the standard Reveal policy in NS_Reveal ($p$ = 0.029). [^19] Appendix Figure A1 illustrates the match rate across rounds for each treatment, suggesting that the observed differences persist throughout the experiment. These findings are consistent with Hypothesis 1.
@@
To summarize participants’ adherence to equilibrium strategies across treatments, we construct a statistic termed average absolute deviation (AAD), which calculates the treatment-level frequency of deviating from the equilibrium strategies summarized in Table 2. 2(b) displays the results, showing that the Reveal policy substantially reduces AAD irrespective of the presence of neighborhood schools. In the absence of neighborhood schools, the Reveal policy reduces AAD from 0.42 to 0.04 ($p$ = 0.029), meaning that only 4% of decisions deviate from equilibrium strategies. When neighborhood schools are available, the Reveal policy reduces AAD from 0.39 to 0.21, and the RevealMore policy further reduces it to 0.07 ($p$ = 0.029 in all treatment comparisons). 2(b) further shows that treatment differences persist across rounds, although AAD generally decreases over time, reflecting learning effects; this reduction is especially pronounced in NS_Reveal. [^20] Appendix Table A1 reports complementary regression analyses confirming that treatment differences are significant both across all rounds and during the last 5 rounds, with a noticeable decline in suboptimal decisions in NS_Reveal during the last 5 rounds.
@@
Note that in NS_Reveal, no-neighborhood students’ strategies are “non-monotonic” in their lottery numbers. Specifically, their equilibrium strategy is to report $s_{1}$ when the lottery number is 5, which is unintuitive. Indeed, experimental data indicate that participants deviated from this strategy over 95% of the time under this circumstance (see Table 5). Therefore, we performed a robustness check by excluding observations for no-neighborhood students with lottery number 5 across all treatments (to ensure fair comparisons). AAD in NS_Reveal decreases to 0.14, which remains significantly higher than that in NS_RevealMore (pp=0.029). However, Appendix 2(b) suggests that during the last 5 rounds, AAD levels between NS_Reveal and NS_RevealMore become nearly indistinguishable, a finding corroborated by complementary regression analyses reported in Appendix Table A2. This finding suggests that the efficiency gain associated with providing the additional lottery information tends to diminish as participants gain experience. This is nonetheless consistent with our earlier discussion at the end of Subsection 2.2, which posits that increased market stability may reduce the necessity for full revelation of lottery information. Overall, the analysis of AAD offers initial empirical support for Hypothesis 3.
@@
Table 5 shows that, in the absence of neighborhood schools, the Reveal policy significantly increases the likelihood that students with various lottery numbers adhere to equilibrium strategies. When neighborhood schools are available, the Reveal policy mainly benefits $s_{3}$-neighbor and no-neighborhood students by increasing their conformity to equilibrium strategies. [^21] The Reveal policy also slightly reduces the overall frequency of deviation for $s_{2}$-neighbor students from 8.75% to 6.25%. However, $s_{2}$-neighbor students with lottery number 2 tend to make more mistakes than those with other lottery numbers, occasionally reporting $s_{1}$ despite an equilibrium strategy to always report $s_{2}$. For example, $s_{3}$-neighbor students deviate only 15% of the time, which is substantially lower than 68.13% in NS_Cover. [^22] In previous literature, this behavioral pattern of $s_{3}$-neighbor students observed in NS_Cover is referred to as neighborhood school bias (calsamiglia2010constrained). However, given that theoretical predictions vary across policies, we refrain from asserting that the Reveal policy directly mitigates this bias here. Notably, the efficacy of the Reveal policy is not uniform across lottery numbers; students receiving either the best or worst lottery numbers deviate much less frequently than those with intermediate lottery numbers. As previously noted, the equilibrium strategy for no-neighborhood schools with lottery number 5 is unintuitive, causing almost all these students’ decisions to deviate from the equilibrium strategy. However, even after excluding these observations, the tendency of $s_{3}$-neighbor and no-neighborhood students with intermediate lottery numbers making more mistakes remains persistent. To further investigate the sources of these mistakes, Table 6 presents the frequencies of all possible submission strategies by students’ roles and utility types. Compared to theoretical predictions summarized in Table 3, $s_{3}$-neighbor and no-neighborhood students tend to err on the side of being too conservative in their submission strategies under the Cover policy. Although the Reveal policy mitigates this tendency, particularly among type-$\mathbf{v}$ students, it does not entirely eliminate it. This behavioral pattern suggests that $s_{3}$-neighbor and no-neighborhood students with intermediate lottery numbers face more complex strategic challenges, which the Reveal policy partially alleviates, relative to $s_{1}$-neighbor, $s_{2}$-neighbor and students with either the best or worst lottery numbers.
@@
Finally, panel (3) of Table 5 shows that the RevealMore policy further enhances adherence to equilibrium strategies among $s_{3}$-neighbor and no-neighborhood students, notwithstanding the persistence of relatively higher deviation rates among those with intermediate lottery numbers. For example, $s_{3}$-neighbor students deviate only 6.25% of the time; even those with lottery number 3 err in just 14.29% of cases, substantially lower than 46.43% in NS_Reveal. Appendix Table A3 reports more detailed analyses, allowing decisions to depend on both own lottery number and lottery numbers of neighborhood students. We observe that except in a few instances with limited observations (i.e., no more than 6), the frequencies of deviation generally remain below 20%. One notable exception occurs when no-neighborhood students with lottery number 4 observe that both $s_{1}$-neighbor’s and $s_{2}$-neighbor’s lottery numbers exceed 2. In this case, although they should have reported $s_{3}$, they mistakenly report $s_{2}$ in 35% of the time. This discrepancy likely reflects the strategic complexity inherent in the situation: despite holding a better-than-average lottery number of 3, which intuitively suggests a favorable chance of admission to $s_{2}$, the equilibrium strategy indicates otherwise.
@@ 3.3 A Robustness Experiment in More Complex Environments
Our main experiment examines scenarios in which students are permitted to report only one school in their preference list. The matching market is relatively small and the cardinal preferences are somewhat arbitrary. However, this setup is essential for making precise theoretical predictions, but perhaps at the cost of realism. To partially address this limitation, the robustness experiment considers a larger matching market with neighborhood schools (16 students and 4 schools), allowing students to report up to two schools. All students’ ordinal preferences are identical, but their cardinal preferences incorporate multiple random components. The principal findings are consistent across both experiments: the Reveal policy improves the overall match rate and market efficiency compared to the Cover policy, and students generally make more informed decisions based on observed lottery numbers. A detailed report on the robustness experiment is provided in Online Appendix B.
@@ 4 Conclusion
Lotteries are widely used in school choice systems to break priority ties. This paper advocates for a policy that reveals lottery results to students before preference submission, contrasting it with the prevailing practice of withholding such information. Our theoretical analysis shows that when students face constraints on the length of their preference lists, which is a common feature in practice, the reveal policy improves students’ decision-making, enabling them to coordinate their choices more effectively. This, in turn, enhances matching outcomes and increases overall welfare. Complementary lab experiments support these findings, showing that the reveal policy leads to higher match rates, higher student payoffs, and better alignment between reported preferences and achievable options.
@@
Appendix: Additional Figures and Tables
@@
Online Appendix
Appendix A Experimental Instructions
@@
Appendix B Robustness Experiment
B.1 Experimental Design
@@
The experiment was conducted at the Nanjing Audit University Economics Experimental Lab with a total of 192 university students, using the software z-Tree (Fischbacher2007). We ran six sessions for each treatment. Each session consisted of 16 participants who interacted for a total of 40 rounds, divided into two blocks of 20 rounds each. Within each group, each participant was randomly assigned a student ID (ranging from 1 to 16), a lottery number (also ranging from 1 to 16), and payoffs from attending each of the four schools; all of these values were distinct in each round. Specifically, we generated a sequence of the 6-tuple of ID, lottery number and four payoffs at the group-round level. We applied it to each session to alleviate the concern that any treatment difference is simply due to the different random numbers. After every round, all participants received feedback about whether and which school they were admitted to and their round payoff. At the end of the session, one round from each block was privately and randomly chosen for each participant as her payoff for that block, and the participant’s total payoff was the sum of the payoffs from the two blocks. The experimental instructions are reproduced in the online appendix.
@@ B.2 Experimental Results
Figure B1 shows the match rate across rounds under each condition. Consistent with our hypothesis, the reveal policy results in a significantly higher match rate than the cover policy, irrespective of the ROL length. When the ROL length is one, the average match rate is 89.1% under the reveal policy, which is significantly higher than 77.8% under the cover policy ($p$ = 0.004, Wilcoxon rank-sum test with each session considered as an independent observation). Similarly, when the ROL length is two, the rate is 98.5% under the reveal policy compared to 85.9% under the cover policy ($p$ = 0.003).
Figure B1: Match rate over round
@@
The reveal policy is expected to assist students in making more rational decisions and thereby reduce their propensity to play safe by frequently reporting neighborhood schools. Table B1 presents the neighborhood school report rate under each condition, focusing on students who have neighborhood schools. The table also presents a breakdown of this statistic by each of the four lottery categories. For instance, lottery (1, 4) refers to the four students whose lottery numbers are 1, 2, 3, and 4.
@@
Table B1: Neighborhood school report rate and bias
@@
The reveal policy is expected to reduce not only the neighborhood school bias in students’ ROL submissions but also the segregation level in the final admission outcome, measured as the proportion of students assigned to their neighborhood schools. Table B2 in Online Appendix Appendix: Additional Figures and Tables indicates that the reveal policy results in a significantly lower segregation level compared to the cover policy, but this effect is significant only when the ROL length is one. When the ROL length is two, although the effect aligns with expectations, the reveal policy exhibits a weaker and statistically insignificant effect on the segregation level. This is primarily driven by the luckiest students who could still be admitted to schools that are better than their neighborhood schools when they have more than one school to report. Thus, their neighborhood school bias has a much weaker impact on their final admission outcomes.
Table B2: Segregation (the proportion of subjects assigned to their neighborhood schools, only for those who have a neighborhood school)
@@
Table B3 reports the efficiency, measured by the student’s average payoff, under each condition. Consistent with our hypothesis, the reveal policy results in significantly greater overall efficiency than the cover policy, irrespective of the ROL length. However, the benefits of the reveal policy are not uniformly distributed among students. While students with the best or worst lottery numbers tend to benefit from this policy, students with intermediate lottery numbers tend to suffer. The issue of unevenly distributed benefits is pronounced when the ROL length is one, but it is largely mitigated when the ROL length is two. Intuitively, when the possibility of mismatch is high (i.e., when the ROL length is short), prior knowledge of lottery numbers assists the most fortunate students in gaining admission to better schools while helping the least fortunate students avoid being left unmatched. However, this dynamic tends to disadvantage students in the middle, making them less likely to gain admission to top schools.
Table B3: Efficiency (mean payoff)
@@
Following our review of the aggregate-level results, we turn to individual ROL strategies. Table B4 presents the frequency of each possible strategy under each condition. Under the reveal policy, we also calculate the frequency for each lottery category separately. When the ROL length is one, students tend to report better schools more frequently under the cover policy; in contrast, under the reveal policy, the frequency of reporting each school is relatively uniform. The reason is that students make more informed decisions under the reveal policy by primarily reporting schools where they have a realistic chance of admission: students whose lottery numbers range from 1 to 4 are most likely to report school A; students whose lottery numbers range from 5 to 8 are most likely to report school B; and so on. Similarly, when the ROL length is two, under the cover policy, students tend to report a pair of schools that more frequently include at least one top school (such as A-B and A-C) than two bottom schools (such as C-D); conversely, under the reveal policy, the reported pairs of schools are evenly distributed between top and bottom schools (such as A-B, B-C and C-D). Further, students tend to report schools to which they have realistic chances of admission based on their lottery numbers: students whose lottery numbers range from 1 to 4 are most likely to report the school pair A-B; students whose lottery numbers range from 5 to 8 are most likely to report the school pair B-C; students whose lottery numbers range from 9 to 16 are most likely to report the school pair C-D. [^23] We also calculate the frequency of each possible strategy for students who have a specific neighborhood school (Table B5, Table B6 and Table B7) and for students who do not have any neighborhood school (Table B8) separately. The results generally align with our expectations. For instance, students who have neighborhood school B only report school A or B when the ROL length is one and the school pairs A-B, B-C or B-D when the ROL length is two.
Table B4: Proportion of each possible submission strategy (%)
@@
Table B5: Proportion of each possible submission strategy (%) for students having neighborhood school B
@@
Table B6: Proportion of each possible submission strategy (%) for students having neighborhood school C
@@
Table B7: Proportion of each possible submission strategy (%) for students having neighborhood school D
@@
Table B8: Proportion of each possible submission strategy (%) for non-neighborhood students